OpenStudy (anonymous):

Find the length of the radius of a sphere with a surface area of 64.62 km2.

OpenStudy (anonymous):

Sphere's surface area is 4*pi*r^2 = 64.62km2. r = radius of sphere

OpenStudy (anonymous):

Ööh, do you know how to solve that equation for r? :)

OpenStudy (mertsj):

\[SA=4\pi r^2\]

OpenStudy (mertsj):

To solve that for r first divide both sides by 4pi

OpenStudy (mertsj):

\[\frac{SA}{4\pi }= r^2\]

OpenStudy (mertsj):

Now take the square root of both sides to find r

OpenStudy (mertsj):

\[\sqrt{\frac{SA}{4\pi }}=r\]

OpenStudy (mertsj):

Now replace SA with 64.62 and do the arithmetic.

OpenStudy (anonymous):

So, I got 2.27 km. Is this correct? (:

OpenStudy (mertsj):

yes